Cori10:

A standard showerhead in Marty's house dispenses 7 gallons of water per minute. Marty changed this showerhead to an energy-saving one. The equation shows the amount of water dispensed, y, as a function of the number of minutes, x, for the new showerhead. y = 3x How much water does Marty save each day with the change in showerheads if he uses the shower for 8 minutes each day? 4 gallons 12 gallons 32 gallons 53 gallons

supie:

It was given that a showerhead dispenses \[7\ gallons\ of\ water/min\] We have to find If he showers for 8 minutes, he uses how much water? We know that \[1\ minute = 7\ gallons\] and \[8\ minutes = x\] so \[x = 7 × 8\ minutes = 56\ gallons\ of\ water\] The problem gave us that "hows the amount of water dispensed, y, as a function of the number of minutes, x, for the new showerhead. y = 3x" The equation that needs to be used for saving energy in the shower\[y = 3x\] \[(x=number\ of\ minutes\ spent\ in\ the\ shower)\] \[8\ minutes\]\[y = 3 × 8\]\[y = 24\ gallons\] He uses \[8\ minutes\ /\ day\] Which is going to be computed as \[56\ gallons - 24\ gallons\]\[which = 32 gallons\]